# Inertial Torsion Noise in Matter-Wave Interferometers for Gravity Experiments

### v1

@@ VII Summary
In section III, we pointed out that the decay factor $\Gamma$ can be regarded as a linear response of the interferometer to the ITN. In particular, it can be formulated as the power spectrum density of the noise $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ multiplying a transfer function $F(\omega)$ in frequency space. The transfer function only relies on the trajectories of the interferometer’s two arms and is independent of the noise. The exact and approximate formulae of $F(\omega)$ are given by Eqs. (13) and (14), respectively.
In section IV, we considered a specific ITN source: the thermal gas molecule collisions on the experiment box. The motion of the box is determined by the Langevin equation, by which one can compute the PSD $S_{\Theta\Theta}(\omega)$ described in Eq. (21). According to the convolution theorem, the PSD of the ITN $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ is the self-convolution of $S_{\dot{\Theta}\dot{\Theta}}(\omega)$ which relates to $S_{\Theta\Theta}(\omega)$ through Eq. (16). The calculation details of $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ are summarized in Appendix B. We also did a Monte-Carlo simulation directly from the dynamical equation in Eq. (18) to double-check the analytical result. Another essential feature of ITN $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ is that the peak position is doubled compared to $S_{\dot{\Theta}\dot{\Theta}}(\omega)$ because of properties of self-convolution. In addition, we showed that the Q-factor retains the usual expression as another safety check.
Finally, we studied the dephasing factor caused by ITN in section V. The dephasing factor increases significantly near the resonance region $\Omega_{\rm rot}=2n\pi/T_{\rm tot}$ for some integer $n$. It increases proportional to the damping rate $\gamma_{\rm rot}$ in the underdamped region, while it decreases with $\gamma_{\rm rot}$ in the overdamped region. As is shown in Fig. 5 and Fig. 6, the maximum value of $\Gamma$ is tiny so the dephasing caused by the thermal gas molecules collision is negligible.
For a generic source of noise described by a generalized Langevin equation (28), the dephasing factor is proportional to $\gamma_{\rm rot}^{-1}$ and $\gamma_{\rm rot}^{-3}$ in the underdamped and overdamped regions, respectively. In Fig. 9 we computed the constraints on the experimental parameters such that the dephasing factor remains smaller than $0.01$. We found that the amplitude of the noise has to be smaller than $10^{-10}\,\rm Hz^{3}$ in the resonance region, while the constrain can be relaxed to $10^{-5}\,\rm Hz^{3}$ outside the resonance region (with a damping rate larger than $10^{-4}$ Hz). Finally, we briefly compared our results with the rotational noise analysis in the context of cold atom interferometer on-board a satellite [52].
In summary, this work quantifies the dephasing in nanoparticle matter-wave interferometry arising from inertial torsion noise (ITN). We developed the general methodology as well as obtained constrains for specific sources of noise for a simple single-stage suspension forming a torsion pendulum. In future it would be interesting to perform the noise analysis for more refined suspensions such as the inverted pendulum [42] and the Roberts linkage [43].
@@ Acknowledgements.
M. Wu would like to thank the China Scholarship Council (CSC) for financial support. MT would like to acknowledge funding from ST/W006227/1. SB thanks EPSRC grants EP/R029075/1, EP/X009467/1, and ST/W006227/1. SB and AM’s research is supported by the Sloane, and Betty and Moore foundations.
@@ A.2 Lagrangian of Inertial Torsion Noise
Te Lagrangian of a point-like massive object is given by $L=-mc^{2}\sqrt{-\mathrm{d}s^{2}/(c^{2}\mathrm{d}t^{2})}$. Using the metric in Eq. (36) and taking the non-relativistic limit $v_{x},v_{y}\ll c$ (with $v_{x}\equiv\mathrm{d}x/\mathrm{d}t,v_{y}\equiv\mathrm{d}y/\mathrm{d}t$), we find that the Lagrangian is

### v2

@@ VI Summary
In section III, we pointed out that the decay factor $\Gamma$ can be regarded as a linear response of the interferometer to the ITN. In particular, $\Gamma$ can be formulated as the PSD of the noise $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ multiplying a transfer function $F(\omega)$ in frequency space. Remarkably, $F(\omega)$ only relies on the trajectories of the two arms of the interferometer and is independent of the noise. In the rest part of section 3, some asymptotic features of $F(\omega)$ are discussed, and the exact and approximate result for a certain interferometer is shown as Fig. 2.
In section IV, we modelled the torsion noise by a generalized Langevin equation (17), which implies the PSD $S_{\Theta\Theta}(\omega)$ as (19). Then the PSD of the ITN $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ is the self-convolution of $S_{\dot{\Theta}\dot{\Theta}}(\omega)$ according to the convolution theorem, where the math details are summarized in Appendix B. It is remarkable that the peak position of $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ is doubled compared to $S_{\dot{\Theta}\dot{\Theta}}(\omega)$ because of properties of self-convolution, while the Q-factor of $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ remains the same as $S_{\dot{\Theta}\dot{\Theta}}(\omega)$.
In section V, we scanned parameters theoretically and found some major features as follows. First, $\Gamma$ increases significantly near the resonance region $\Omega_{\rm rot}=2n\pi/T_{\rm tot}$. Next, $\Gamma$ descreases proportional to the damping rate $\gamma^{-1}$ in the underdamped region, while it decreases with $\gamma^{-3}$ in the overdamped region. Finally, if $\Gamma$ is required smaller than $0.01$, then constrain on the parameters $A$, $\gamma$ and $\Omega_{\rm rot}$ shown as Fig.8, and $A$ is tolerant up to $10^{-5}\,\rm Hz^{3}$ if $\gamma>10^{-4}$ Hz and $\Omega_{\rm rot}$ is outside the resonance region.
@@ Acknowledgements.
M. Wu would like to thank the China Scholarship Council (CSC) for financial support. MT would like to acknowledge funding from ST/W006227/1. SB thanks EPSRC grants EP/R029075/1, EP/X009467/1, and ST/W006227/1. SB and AM’s research is supported by the Sloan, and Betty and Moore foundations.
@@ A.2 Lagrangian of Inertial Torsion Noise
The Lagrangian of a point-like massive object is given by $L=-mc^{2}\sqrt{-\mathrm{d}s^{2}/(c^{2}\mathrm{d}t^{2})}$. Using the metric in Eq. (31) and taking the non-relativistic limit $v_{x},v_{y}\ll c$ (with $v_{x}\equiv\mathrm{d}x/\mathrm{d}t,v_{y}\equiv\mathrm{d}y/\mathrm{d}t$), we find that the Lagrangian is
@@ Appendix B Calculation for PSD of ITN
Appendix C Gas collision noise
In this appendix, we will consider a certain source of the inertial torsion noise, that is, the collision due to thermal motion of gas molecules. According to the fluctuation-dissipation theorem, the applitute of the random force due to the gas collision is $\sqrt{A}=\sqrt{2\gamma_{\rm rot}k_{B}T/I}$, where $k_{B}$ is the Boltzmann constant and $T=300$ K is the gas temperature outside the experiment box. Then the dynamical equation (17) of the torsion motion of the experiment box becomes to the Langevin equation [64]
In this case, the power spectrum for $\Theta$ and $\dot{\Theta}^{2}$ are
$S_{\Theta\Theta}(\omega)$ $=\frac{2\gamma_{\rm rot}k_{B}T/I}{(\Omega_{\rm rot}^{2}-\omega^{2})^{2}+\gamma_{\rm rot}^{2}\omega^{2}},$ (40)
$S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ $=\frac{4\pi\gamma_{\rm rot}(k_{B}T)^{2}}{I^{2}}\frac{4\omega^{4}+4(\gamma_{\rm rot}^{2}-3\Omega_{\rm rot}^{2})\omega^{2}+16\Omega_{\rm rot}^{4}}{(\omega^{2}+\gamma_{\rm rot}^{2})(4\gamma_{\rm rot}^{2}\omega^{2}+(\omega^{2}-4\Omega_{\rm rot}^{2})^{2})}.$
It is notatble that $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ has an additional $\gamma^{2}$ dependence in comparison to the general case in Eq. (20). This difference arises because the amplitude $A=2\gamma_{\rm rot}k_{B}T/I$ of the external force for the thermal noise in Eq. (39) is proportional to the damping rate $\gamma_{\rm rot}$, which does not hold for the general case considered in this section.
Note that in this case, the damping rate $\gamma_{\rm rot}$ describes both the dissipation effect and the random force caused by the collision from the ambient thermal gas molecules, which is given by [65, 66]
where $P_{\rm gas}$ is the pressure of gas and $m_{\rm gas}$ is the mass of the gas molecules. Since $\gamma_{\rm rot}$ is proportional to $P_{\rm gas}$, it can vary depending on the gas pressure outside the box. Apart from $P_{\rm gas}$, all other factors contribute a factor around $10^{-4}\sim 10^{-3}$ Hz/Pa, so the value of the damping rate $\gamma_{\rm rot}$ can be estimated as
$\gamma_{\rm rot}/[\rm Hz]\sim 10^{-4}P_{\rm gas}/[\rm Pa].$ (42)
Figure 11: Dephasing factor $\Gamma$ as a function of the damping factor $\gamma_{\rm rot}$ for the thermal case. The transfer function is given in Eq. (11) with parameters $t_{a}=0.25$ s and $t_{e}=0$ s. Three different $\Omega_{\rm rot}$ are chosen in the low rotation frequency limit (blue line), the resonance frequency condition $\Omega_{\rm rot}\sim 2\pi/T=2\pi$ Hz (orange line), and the high rotation frequency limit (green line). The dephasing $\Gamma$ is proportional to $\gamma_{\rm rot}$ when $\gamma_{\rm rot}\ll\Omega_{\rm rot}$, while it starts decreasing with respect to $\gamma_{\rm rot}$ when $\gamma_{\rm rot}$ becomes comparable or larger than $\Omega_{\rm rot}$.
As for the dephasing factor, Fig. 11 shows the dependence of $\Gamma$ on the damping factor $\gamma_{\rm rot}$ caused by the thermal gas molecules. As is shown, for the thermal case, $\Gamma$ is proportional to $\gamma_{\rm rot}$ and $\gamma_{\rm rot}^{-1}$ in the underdamped region and the overdamped region in respect, which has an additional $\gamma^{2}$ dependence in comparison to the general case because the thermal amplitude $A^{2}$ is proportional to $\gamma_{\rm rot}^{2}$.
A final remark on Fig.11 is that the dephasing factor $\Gamma$ caused by the torsion noise from gas molecules collision is negligibly tiny. Two main reasons cause this. First, the thermal motion of gas molecules is proportional to a small thermal factor $k_{B}T\sim 10^{-21}$ J. Besides, ITN is a second-order effect, so every little factor in the PSD $S_{\Theta\Theta}(\omega)$ of the experiment apparatus will be squared for the dephasing factor $\Gamma$. Combining both effects, the dephasing factor $\Gamma$ is suppressed by a highly tiny factor $(k_{B}T)^{2}\sim 10^{-42}\,\rm J^{2}$, such that $\Gamma$ does not exceed $10^{-30}$.